OpenStudy (anonymous):

Karen and Jeremy are purchasing a house and are financing $225,000. The mortgage is a 30-year 4/1 ARM at 3.0% with a 2/6 cap structure. What will the remaining balance be after the first 4 years? $196,208.00 $210,667.50 $205,335.00 $210,895.73

OpenStudy (amistre64):

its a 4year initial term, so define the payments and run a balance

OpenStudy (amistre64):

P = B4 = Bk^(4*12)-P(1-k^(4*12))/(1-k) Bk^(4*12)-(Bk^(30*12)(1-k)/(1-k^(30*12)))(1-k^(4*12))/(1-k) 225000k^(4*12)-(225000k^(30*12)(1-k)/(1-k^(30*12)))(1-k^(4*12))/(1-k), k=1+.03/12 http://www.wolframalpha.com/input/?i=225000k%5E%284*12%29-%28225000k%5E%2830*12%29%281-k%29%2F%281-k%5E%2830*12%29%29%29%281-k%5E%284*12%29%29%2F%281-k%29%2C+k%3D1%2B.03%2F12

OpenStudy (amistre64):

looks like a cut instead of copied my P lol

OpenStudy (amistre64):

P = Bk^(30*12)(1-k)/(1-k^(30*12))